# Efficient Low Rank Attention for Long-Context Inference in Large Language Models

### v2

@@ 5 Experiments
In this section, we conduct a series of experiments to rigorously evaluate the effectiveness and performance of our proposed method. The experiments are performed on NVIDIA A100 GPUs. The number of max iteration and the tolerance introduced in Algorithm 1 and 2 are chosen as 2 and 0.01, respectively. All scaling parameters are set as $\lambda_{pQ}=\lambda_{pK}=\lambda_{d1}=\lambda_{d2}=1$. All LLM evaluations are empowered by OpenCompass [27]. Additional results can be found in Appendix B, the time cost analysis is presented in Appendix C, and a guideline for hyperparameter tuning is provided in Appendix D.
@@ 5.1 Accuracy Evaluation
We evaluate the accuracy of the proposed method on the RULER dataset [28], using a long-context setting with a sequence length of 128K tokens. For the low rank approximation, the rank is set to $r=32$. The number of top-$k$ tokens selected based on attention scores is set to 2048 (1.56% of 128K), while the number of lite tokens, i.e., the most recently generated tokens, is set to 64. The backbone language model used in the evaluation is LLaMA-3-8B-1M [24]. We compare our method against four recent dynamic sparse attention baselines: Loki [21], InfiniGen [14], Quest [20], and ShadowKV [16]. Results are shown in Table 1.
@@
We further evaluate the proposed method on two backbone models: LLaMA-3-8B-1M [24] and Qwen2.5-7B-Instruct [2], using the same configuration of rank $r=16$, top-$k=256$ (6.25% of 4K), and 16 lite tokens on a subset of RULER-4K. As shown in Table 2, our method consists the performance of both models.
@@ 5.2 Impact of Rank and Top-$k$ Selection
In this subsection, we investigate the impact of two key hyper-parameters in our method: the rank $r$ used in the low rank approximation, and the number of top-$k$ active tokens selected based on attention scores.
Rank Selection. To isolate the effect of the rank, we fix the number of top-$k$ tokens at 256 and the number of lite tokens at 16. We vary the rank parameter over the set $\{16,24,32\}$, allowing us to examine how the capacity of the low rank representations affects model performance.
@@ 5.3 Miss Rates
We evaluate the performance of the proposed strategy for memory management by computing the miss rate. The rate is a ratio $c_{\text{miss}}/c_{\text{total}}$, where $c_{\text{miss}}$ denotes the number of KV cache rows that must be transferred from CPU to GPU, and $c_{\text{total}}$ is the number of all selected indices. This metric quantifies the efficiency of the token selection mechanism in reducing memory transfers.
Experiments are conducted on the summarization task using the wikitext-2-v1 test set [25]. We perform a grid search over different hyper-parameters, including the rank $r\in\{8,16,32,64\}$, the number of active tokens $\in\{128,256,512\}$, and the number of lite tokens $\in\{4,8\}$. The distribution of miss rates is illustrated in Figure 5, which reveals an approximately Gaussian shape with a mean miss rate of around 0.40 and a standard deviation of approximately 0.10. This implies that, on average, the proposed hit-and-miss mechanism reduces the amount of data transferred from CPU to GPU by about 60% in average.
@@ 6 Conclusion
While LRQK significantly reduces data transfer overhead, further observation reveals that CPU-side indexing operations constitute the primary performance bottleneck rather than PCIe bandwidth limitations, and hoping can be addressed in future work. Additionally, LRQK’s hyper-parameters require task-specific tuning. While we provide practical guidelines and default configurations, optimal settings still demand empirical validation for new domains.
@@ A.2 Decode Derivatives
To update $\mathbf{B}_{Q,t-1}$ and $\mathbf{B}_{K,t-1}$, we utilize gradient descent. The gradient of $\mathcal{L}_{\text{dec}}$ with respect to $\mathbf{B}_{Q,t-1}$ and $\mathbf{B}_{K,t-1}$ are:
@@
Similarly, for $\mathbf{B}_{K,t-1}$, we can compute the optimal $\eta_{K}^{*}$ as,
@@ B.4 Results on Larger Models
To evaluate the scalability of LRQK to larger models, additional experiments are conducted on ‘Qwen/Qwen2.5-14B-Instruct’(Qwen 14B) with 64K context length and ‘Qwen/Qwen2.5-32B-Instruct’ (Qwen 32B) with 16K context length under A100 80G GPU. Following the official guidance from Qwen2.5, we apply YaRN [29] for positional encoding extension in the 64K context setting. All experiments use the default LRQK hyperparameters (rank=32, top-2048, and 64 lite tokens).
@@ B.5 Results With KVQuant
To investigate whether LRQK is compatible with quantization-based compression methods, we evaluate the combination of LRQK with KVQuant [6], a representative KV cache quantization approach. We conduct experiments on ‘meta-llama/Llama-3.1-8B-Instruct’ (Llama-3.1-8B) and ‘Qwen/Qwen2.5-7B-Instruct’ (Qwen 2.5-7B-Instruct) using the RULER 32K benchmark. The LRQK hyperparameters are set to $r=32$, top-$k=2048$, and 64 lite tokens.
@@
Table 7: Performance comparison on RULER 32K with KVQuant applied alone and in combination with LRQK. Columns ‘+KVQuant‘ indicate KVQuant only, and columns ‘++LRQK‘ indicate both KVQuant and LRQK applied.
@@ B.6 Impact of Initialization Strategies
The second strategy aims to minimize reconstruction error by selecting the most salient dimensions from $\mathbf{Q}$ and $\mathbf{K}$ independently. We compute the $L_{1}$ norm of each dimension across the sequence:
@@
These three initialization strategies are evaluated on ‘meta-llama/Llama-3.1-8B-Instruct-1M‘ and ‘Qwen/Qwen2.5-7B-Instruct‘ using the RULER 16K benchmark with default LRQK hyperparameters ($r=32$, top-$k=2048$, 64 lite tokens).
@@ C.1 Throughput Results
We evaluate the runtime performance of LRQK on ‘meta-llama/Llama-3.1-8B-Instruct-1M’, comparing it against standard GPU-only and CPU offloading approaches. Experiments are conducted using the Hugging Face transformers library [^1] https://huggingface.co/docs/transformers/v4.52.2/en/main_classes/text_generation on a single NVIDIA A100 GPU (40 GB memory) with batch size 1. We use the RULER QA-2 (HotpotQA) dataset with context lengths ranging from 4K to 64K tokens. The parameters of LRQK are set as $r=16$, top-$k=1024$ active tokens, and 16 lite tokens.
@@
CPU offload enables processing of longer contexts but suffers severe throughput degradation, particularly during decoding at 32K–64K tokens. This performance drop stems from the need to transfer the entire KV cache between CPU and GPU for each decoding step, creating a bottleneck that scales linearly with context length.
@@ C.2 Comparison with Baseline Methods
To provide a more comprehensive performance analysis, we compare LRQK against vanilla attention and ShadowKV[16] on a text summarization task using 20 samples from LongBench with 32K context and up to 128 output tokens. Method ShadowKV is configured with its default parameters (sparse budget = 2048, rank=160, chunk size=8), and the definition of ‘rank’ is not the same as this paper, please refer to the original paper for more details. Experiments are conducted on NVIDIA GeForce RTX 3090 (24 GB) GPUs. Due to memory constraints, vanilla attention requires 2 GPUs, while ShadowKV and LRQK operate on a single GPU.
@@ Appendix D Hyperparameter Selection Guidelines
LRQK introduces several hyper-parameters that require configuration: rank $r$ for low rank approximation, top-$k$ for active token selection, number of lite tokens, and convergence parameters (iterations and tolerance). While optimal values vary across model architectures and tasks, we provide practical guidelines based on our extensive experiments to minimize tuning effort.
@@
We recommend starting with the following default configuration, which achieves strong performance across diverse settings: rank $r=32$, active tokens top-$k=2048$, lite tokens $64$, iterations $2$, and tolerance $10^{-2}$. This configuration provides a balanced trade-off between approximation quality, memory efficiency, and computational overhead, and serves as an effective starting point for task-specific tuning.
@@
The rank $r$ controls the dimensionality of the low rank factors $\mathbf{A}_{Q},\mathbf{A}_{K}\in\mathbb{R}^{\cdots l\times r}$ for each attention head independently. Since typical attention heads have dimension $d_{\text{head}}=128$, the theoretical range for $r$ is $[1,128]$. In practice, we find that $r\in\{8,16,32,48\}$ can provide a favorable balance between approximation quality and computational cost. Lower ranks (e.g., $r=8$) reduce memory usage but may sacrifice accuracy, while higher ranks (e.g., $r=48$) improve approximation at the cost of increased computation. For most applications, $r=32$ offers an effective compromise.
@@
The number of lite tokens specifies how many recent tokens are always retained in GPU memory to capture local context. Recommend choices are $\{16,32,64\}$, with minimal impact on overall computational cost. This parameter is relatively insensitive; we recommend $64$ lite tokens as a conservative default that ensures sufficient local context coverage.
@@
The search space for LRQK hyper-parameters is relatively compact compared to other optimization-based methods. We recommend the following tuning protocol:

### v3

@@ 5 Experiments
In this section, a series of experiments are conducted to rigorously evaluate the effectiveness and performance of our proposed method. The experiments are performed on NVIDIA A100 GPUs. The number of max iteration and the tolerance introduced in Algorithm 1 and 2 are chosen as 2 and 0.01, respectively. All scaling parameters are set as $\lambda_{pQ}=\lambda_{pK}=\lambda_{d1}=\lambda_{d2}=1$. All LLM evaluations are empowered by OpenCompass [27]. Additional results can be found in Appendix B, the time cost analysis is presented in Appendix C, and a guideline for hyperparameter tuning is provided in Appendix D.
@@ 5.1 Accuracy Evaluation
The accuracy of the proposed method is evaluated on the RULER dataset [28], using a long-context setting with a sequence length of 128K tokens. For the low rank approximation, the rank is set to $r=32$. The number of top-$k$ tokens selected based on attention scores is set to 2048 (1.56% of 128K), while the number of lite tokens (the most recently generated tokens) is set to 64. The backbone language model used in the evaluation is LLaMA-3-8B-1M [24]. This method is compared against four recent dynamic sparse attention baselines: Loki [21], InfiniGen [14], Quest [20], and ShadowKV [16]. Results are shown in Table 1.
@@
The proposed method is further evaluated on two backbone models: LLaMA-3-8B-1M [24] and Qwen2.5-7B-Instruct [2], using the same configuration of rank $r=16$, top-$k=256$ (6.25% of 4K), and 16 lite tokens on a subset of RULER-4K. As shown in Table 2, our method consists the performance of both models.
@@ 5.2 Impact of Rank and Top-$k$ Selection
In this subsection, the impact of two key hyperparameters in the proposed method are investigated: the rank $r$ used in the low rank approximation, and the number of top-$k$ active tokens selected based on attention scores.
Rank Selection. To isolate the effect of the rank, we fix the number of top-$k$ tokens at 256 and the number of lite tokens at 16. The rank are vary over the set $\{16,24,32\}$, allowing us to examine how the capacity of the low rank representations affects model performance.
@@ 5.3 Miss Rates
The performance of the proposed strategy for memory management is evaluated by computing the miss rate. The rate is a ratio $c_{\text{miss}}/c_{\text{total}}$, where $c_{\text{miss}}$ denotes the number of KV cache rows that must be transferred from CPU to GPU, and $c_{\text{total}}$ is the number of all selected indices. This metric quantifies the efficiency of the token selection mechanism in reducing memory transfers.
Experiments are conducted on the summarization task using the wikitext-2-v1 test set [25]. A grid search is performed over different hyperparameters, including the rank $r\in\{8,16,32,64\}$, the number of active tokens $\in\{128,256,512\}$, and the number of lite tokens $\in\{4,8\}$. The distribution of miss rates is illustrated in Figure 5, which reveals an approximately Gaussian shape with a mean miss rate of around 0.40 and a standard deviation of approximately 0.10. This implies that, on average, the proposed hit-and-miss mechanism reduces the amount of data transferred from CPU to GPU by about 60% in average.
@@ 6 Conclusion
While LRQK significantly reduces data transfer overhead, further observation reveals that CPU-side indexing operations constitute the primary performance bottleneck rather than PCIe bandwidth limitations, and hoping can be addressed in future work. Additionally, LRQK’s hyperparameters require task-specific tuning. While a practical guidelines and default configurations are provided, optimal settings still demand empirical validation for new domains.
@@ A.2 Decode Derivatives
To update $\mathbf{B}_{Q,t-1}$ and $\mathbf{B}_{K,t-1}$, gradient descent is utilized. The gradient of $\mathcal{L}_{\text{dec}}$ with respect to $\mathbf{B}_{Q,t-1}$ and $\mathbf{B}_{K,t-1}$ are:
@@
Similarly, for $\mathbf{B}_{K,t-1}$, the optimal $\eta_{K}^{*}$ can be computed as,
@@ B.4 Results on Larger Models
To evaluate the scalability of LRQK to larger models, additional experiments are conducted on ‘Qwen/Qwen2.5-14B-Instruct’(Qwen 14B) with 64K context length and ‘Qwen/Qwen2.5-32B-Instruct’ (Qwen 32B) with 16K context length under A100 80G GPU. Following the official guidance from Qwen2.5, YaRN [29] is applied for positional encoding extension in the 64K context setting. All experiments use the default LRQK hyperparameters (rank=32, top-2048, and 64 lite tokens).
@@ B.5 Results With KVQuant
To investigate whether LRQK is compatible with quantization based compression methods, the combination of LRQK with KVQuant [6], a representative KV cache quantization approach, is evaluated. Experiments are conducted on ‘meta-llama/Llama-3.1-8B-Instruct’ (Llama-3.1-8B) and ‘Qwen/Qwen2.5-7B-Instruct’ (Qwen 2.5-7B-Instruct) using the RULER 32K benchmark. The LRQK hyperparameters are set to $r=32$, top-$k=2048$, and 64 lite tokens.
@@
Table 7: Performance comparison on RULER 32K with KVQuant applied alone and in combination with LRQK. Columns ‘+KVQuant’ indicate KVQuant only, and columns ‘++LRQK’ indicate both KVQuant and LRQK applied.
@@ B.6 Impact of Initialization Strategies
The second strategy aims to minimize reconstruction error by selecting the most salient dimensions from $\mathbf{Q}$ and $\mathbf{K}$ independently. The $L_{1}$ norm is computed of each dimension across the sequence:
@@
These three initialization strategies are evaluated on ‘meta-llama/Llama-3.1-8B-Instruct-1M’ and ‘Qwen/Qwen2.5-7B-Instruct’ using the RULER 16K benchmark with default LRQK hyperparameters ($r=32$, top-$k=2048$, 64 lite tokens).
@@ C.1 Throughput Results
Runtime performances of LRQK are evaluated on ‘meta-llama/Llama-3.1-8B-Instruct-1M’, comparing it against standard GPU-only and CPU offloading approaches. Experiments are conducted using the Hugging Face transformers library [^1] https://huggingface.co/docs/transformers/v4.52.2/en/main_classes/text_generation on a single NVIDIA A100 GPU (40 GB memory) with batch size 1. We utilize the RULER QA-2 (HotpotQA) dataset with context lengths ranging from 4K to 64K tokens. The parameters of LRQK are set as $r=16$, top-$k=1024$ active tokens, and 16 lite tokens.
@@
CPU offload enables processing of longer contexts but suffers severe throughput degradation, particularly during decoding at 32K to 64K tokens. This performance drop stems from the need to transfer the entire KV cache between CPU and GPU for each decoding step, creating a bottleneck that scales linearly with context length.
@@ C.2 Comparison with Baseline Methods
To provide a more comprehensive performance analysis, LRQK is compared against vanilla attention and ShadowKV[16] on a text summarization task using 20 samples from LongBench with 32K context and up to 128 output tokens. Method ShadowKV is configured with its default parameters (sparse budget = 2048, rank=160, chunk size=8), and the definition of ‘rank’ is not the same as this paper, please refer to the original paper for more details. Experiments are conducted on NVIDIA GeForce RTX 3090 (24 GB) GPUs. Due to memory constraints, vanilla attention requires 2 GPUs, while ShadowKV and LRQK operate on a single GPU.
@@ Appendix D Hyperparameter Selection Guidelines
LRQK introduces several hyperparameters that require configuration: rank $r$ for low rank approximation, top-$k$ for active token selection, number of lite tokens, and convergence parameters (iterations and tolerance). While optimal values vary across model architectures and tasks, practical guidelines are provided based on our extensive experiments to minimize tuning effort.
@@
It is recommended starting with the following default configuration, which achieves strong performance across diverse settings: rank $r=32$, active tokens top-$k=2048$, lite tokens $64$, iterations $2$, and tolerance $10^{-2}$. This configuration provides a balanced trade-off between approximation quality, memory efficiency, and computational overhead, and serves as an effective starting point for task-specific tuning.
@@
The rank $r$ controls the dimensionality of the low rank factors $\mathbf{A}_{Q},\mathbf{A}_{K}\in\mathbb{R}^{\cdots l\times r}$ for each attention head independently. Since typical attention heads have dimension $d_{\text{head}}=128$, the theoretical range for $r$ is $[1,128]$. In practice, it is found that $r\in\{8,16,32,48\}$ can provide a favorable balance between approximation quality and computational cost. Lower ranks (e.g., $r=8$) reduce memory usage but may sacrifice accuracy, while higher ranks (e.g., $r=48$) improve approximation at the cost of increased computation. For most applications, $r=32$ offers an effective compromise.
@@
The number of lite tokens specifies how many recent tokens are always retained in GPU memory to capture local context. Recommend choices are $\{16,32,64\}$, with minimal impact on overall computational cost. This parameter is relatively insensitive; it is recommended $64$ lite tokens as a conservative default that ensures sufficient local context coverage.
@@
The search space for LRQK hyperparameters is relatively compact compared to other optimization based methods. It is recommended the following tuning protocol: